Can you please help me with this problem
Solve using the square root property
(x + 8)^2 =81
thank you
julie

Answer by stanbon(75887)

Solve using the square root property
(x + 8)^2 =81
---------
x+8 = 9 or x+8 = -9
x = 1 or x = -17
===================
Cheers,
Stan H.